A manufacturing machine has a 1% defect rate.
If 9 items are chosen at random, what is the probability that at least one will have a defect?

Answer by ikleyn(52803)

.

The probability that NO ONE item of 9 has the defect is (1-0.01)^9 = 0.99^9 = 0.913517.


The probability that at least one will have the defect is the COMPLEMENT to it 


    P = 1 - 0.913517 = 0.086483.   (the value under the question).   ANSWER